OpenStudy (anonymous):

The first term in an arithmetic series is 2, while the sum of the first 8 terms of the series is 1472. What is the 8th term in the series? somebody help me

OpenStudy (anonymous):

you got the value of a=2 use the formula S=n/2 [2a+(n-1)d] then calculate d from here.. put it in t=a+(n-1)d

OpenStudy (anonymous):

i dnt kno which number is n d or a

OpenStudy (anonymous):

a = 2 , S8 = 1472 find the d first , 8/2 ( 2(2) + 7d ) =1472 4 ( 4 + 7d ) = 1472 16 + 28d = 1472 d = 52 S8 - S7 = (2 (2) + 7(52) ) - (2 (2) + 6 (52 ) = solve it yourself.. you'll get the answer ;)

OpenStudy (anonymous):

1st term known as 'a' . common difference known as ' d ' :) 'Sn' is sum of terms :)

OpenStudy (anonymous):

*T8 = S8-S7 = (2 (2) + 7(52) ) - (2 (2) + 6 (52 ) = solve it yourself.. you'll get the answer ;)

OpenStudy (anonymous):

im confused on how to solve it

OpenStudy (anonymous):

1st term known as 'a' . common difference known as ' d ' :) 'Sn' is sum of terms :) so , a = 2 , S8 = 1472 find the d first , 8/2 ( 2(2) + 7d ) =1472 4 ( 4 + 7d ) = 1472 16 + 28d = 1472 d = 52 T8 = S8-S7 = ( 2(2) + 7(52) ) - ( 2(2) + 6(52 ) ) = 368 - 316 = 52* 8terms is 52 :)

OpenStudy (anonymous):

*T8 is 8terms :)

OpenStudy (anonymous):

is that help ??

OpenStudy (anonymous):

n=total no. of terms in the sequence,here 8 terms , n=8 a=2...(a is the first term in the sequence) d=?...(which is the common diff.) S=sum of the 8 terms s=n/2 [2a+(n-1)d] 1472=8/2 [2*2 + (8-1)*d] 364=7d d=52 now we know that general term of a sequence is given by T=a+(n-1)d (symbols have same meaning) T=2+(8-1)*52 T=366

OpenStudy (anonymous):

@qiqyn how can the 8th term be 52 if first term is 2 and the common difference is 52?????